OpenStudy (anonymous):

3^3x32+12/4 please help :)

OpenStudy (lgbasallote):

\[3^3 \times 32 + \frac{12} 4 \implies 27 \times 32 + 3\] does that help?

OpenStudy (anonymous):

Yes it does thank you :)

OpenStudy (anonymous):

=867 if you need something to check it against.

OpenStudy (anonymous):

thanks :P